Find the number of 15-centimetre square tiles required to cover a rectangular floor 5.4 m long and 4.05 m wide.
What I have done is :-
5.4 m = 540 cm (since the question wants us to find the number of 15 centimetre square)
4.05 m = 405 cm
Area = 504 x 405 = 204120 cm^2
Since 1 unit of tile = 15 cm^2
Let n be the number of tiles needed to calculate
n x 15 = 204120
n = 204120/15
n=13608
The answer given by my teacher is 972 instead. He always encourages us to try first but this is what I still obtain. I really hope you can show where is my mistake ? Thank you.

Answer by tutorcecilia(2152)

You are on the right tract, but I think you just mixed up 540cm with 504 cm when you did your computations. Just start over again and enter 540cm X 405cm = 218700.
.
Also, try dividing by 15^2 since the tile is a square. Therefore all of its sides are 15cm.
Area = (length)(width) = 15 x 15 = 15^2 = 225:
218700/(15^2) = 972.
Make the correction of n X 15 to n X 15^2 = 218700 and you should get the teacher's answer.